A mother's age is thrice that of her daughter's age. Six years ago, the mother's age was five times that of the daughter's age. What was the daughter's age 6 years ago ?
Answer by Cromlix(4381)

Hi there,
A mother's age is thrice that of her daughter's age.
M = 3D
Six years ago, the mother's age was five times that of the daughter's age.
M - 6 = 5(D - 6)
Substitute 3D from first equation for M
3D - 6 = 5(D - 6)
Multiply out
3D - 6 = 5D - 30
Collect like terms
3D - 5D = -30 + 6
2D = -24
-D = -12
D = 12
Mother's current age = 36
Daughter's current age = 12
Daughter's age 6 years ago = 6 years old
Hope this helps:-)